# Deterministic $(2/3-\varepsilon)$-Approximation of Matroid Intersection

### v1

@@ 1 Introduction
The matroid intersection problem is one of the most fundamental problems in combinatorial optimization. In this problem, we are given two matroids $\mathcal{M}_{1}=(V,\mathcal{I}_{1})$ and $\mathcal{M}_{2}=(V,\mathcal{I}_{2})$ defined on the same ground set $V$ of $n$ elements, and the objective is to find a common independent set $S\in\mathcal{I}_{1}\cap\mathcal{I}_{2}$ of largest possible cardinality, denoted by $r$. This problem generalizes many important combinatorial optimization problems such as bipartite matching, packing spanning trees, and arborescences in directed graphs. Furthermore, it has also several applications outside of traditional combinatorial optimization such as electrical engineering [Mur10, Rec13].
@@
Both the recent $(1-\varepsilon)$-approximation algorithms by Quanrud [Qua24a] and Blikstad–Tu [BT24] are randomized. Therefore, it remains an open question whether a deterministic $\tilde{O}_{\varepsilon}(n)$ independence-query $(1-\varepsilon)$-approximation algorithm can be achieved for almost the entire range of $r$ [^4] Blikstad–Tu [BT24] presented a deterministic $(1-\varepsilon)$-approximation algorithm using a strictly linear number of independence oracle queries only when $r=\Theta(n)$ (and $\varepsilon$ is constant). Then, we consider a deterministic matroid intersection algorithm with only a nearly linear number of independence oracle queries. It is well-known that any maximal common independent set is at least half the size of a maximum common independent set (see e.g., [KV06, Proposition 13.26]). Thus, the natural greedy algorithm—pick an element whenever possible—can be regarded as a deterministic $1/2$-approximation algorithm using a linear number of independence oracle queries. In fact, even beating the trivial $1/2$-approximation ratio for deterministic nearly-linear-independence-query algorithms remains an open problem for almost the entire range of $r$44footnotemark: 4
@@
Given two matroids $\mathcal{M}_{1}=(V,\mathcal{I}_{1})$ and $\mathcal{M}_{2}=(V,\mathcal{I}_{2})$ on the same ground set $V$, for any $\varepsilon>0$, there is a deterministic algorithm that finds a common independent set $S\in\mathcal{I}_{1}\cup\mathcal{I}_{2}$ with $|S|\geq(2/3-\varepsilon)r$, using $O(\frac{n}{\varepsilon}+r\log r)$ independence oracle queries.
@@
Our algorithm uses only a strictly linear number of independence oracle queries when $r=O(n/\log n)$ (and $\varepsilon$ is constant). Prior to our work, Guruganesh–Singla [GS17] presented a randomized $O(n)$-independence-query algorithm that achieves $(1/2+\delta)$-approximation for some small constant $\delta>0$, which was the best approximation ratio for (possibly randomized) algorithms using only a strictly linear number of independence oracle queries.44footnotemark: 4
@@
We note that the time complexity of our algorithm is dominated by the independence oracle queries. That is, our algorithm in Theorem 1 has time complexity $O\left((\frac{n}{\varepsilon}+r\log r)\cdot\mathcal{T}_{\textrm{ind}}\right)$, where $\mathcal{T}_{\textrm{ind}}$ denotes the maximum running time of a single independence oracle query.
@@
Our idea is very simple: we apply a recent $\tilde{O}(n\sqrt{r}/\varepsilon)$-independence-query $(1-\varepsilon)$-approximation algorithm of Blikstad [Bli21], but terminate it before completion. We observe that a $(2/3-\varepsilon)$-approximate solution can be obtained by terminating Blikstad’s algorithm early. As such, our algorithm does not introduce technical novelty, but we believe that it enhances the understanding of algorithms for the matroid intersection problem.
@@
We also implement our algorithm using a rank oracle, which takes as input a subset $S\subseteq V$ and outputs the size of the maximum cardinality independent subset of $S$. Several recent studies have considered fast matroid intersection algorithms in the rank oracle model [LSW15, CLS+19, Tu22]. We note that the rank oracle is more powerful than the independence oracle, since a single query to the rank oracle can determine whether a given set is independent or not.
@@
Chakrabarty–Lee–Sidford–Singla–Wong [CLS+19] presented an $\tilde{O}(n\sqrt{r})$-rank-query exact algorithm and an $O(\frac{n}{\varepsilon}\log n)$-rank-query $(1-\varepsilon)$-approximation algorithm. Their $(1-\varepsilon)$-approximation algorithm requires nearly linear rank oracle queries. Our second result is to obtain a $(2/3-\varepsilon)$-approximation algorithm using a strictly linear number of rank oracle queries.
Theorem 2.
@@
Given two matroids $\mathcal{M}_{1}=(V,\mathcal{I}_{1})$ and $\mathcal{M}_{2}=(V,\mathcal{I}_{2})$ on the same ground set $V$, for any $\varepsilon>0$, there is a deterministic algorithm that finds a common independent set $S\in\mathcal{I}_{1}\cup\mathcal{I}_{2}$ with $|S|\geq(2/3-\varepsilon)r$, using $O(\frac{n}{\varepsilon})$ rank oracle queries.
@@
We note that the time complexity of our algorithm is dominated by the rank oracle queries. That is, our algorithm in Theorem 2 has time complexity $O\left(\frac{n}{\varepsilon}\cdot\mathcal{T}_{\textrm{rank}}\right)$, where $\mathcal{T}_{\textrm{rank}}$ denotes the maximum running time of a single rank oracle query.
@@
Surprisingly, our algorithm can also be implemented in the semi-streaming model of computation. The streaming model for graph problems was initiated by Feigenbaum–Kannan–McGregor–Suri–Zhang [FKM+05], which is called the semi-streaming model. In this model, an algorithm computes over a graph stream using a limited space of $\tilde{O}(n)$ bits, where $n$ is the number of vertices. The maximum matching problem is one of the most studied problems in the graph streaming setting [FKM+05, McG05, EKS09, AG11, GKK12, EKMS12, KMM12, AG13, Kap13, EHM16, KT17, AG18, Kon18, Tir18, GKMS19, ALT21, AB21, AJJ+22, FMU22, HS23, Ber23, KNS23, FS24, Ass24]. The first of these studies was a $(2/3-\varepsilon)$-approximation algorithm for the bipartite maximum matching problem by Feigenbaum et al. [FKM+05], which uses $\tilde{O}(n)$ bits of space and $O(\log(1/\varepsilon)/\varepsilon)$ passes. In particular, whether a $(1-\varepsilon)$ approximation can be achieved in $\text{poly}(1/\varepsilon)$ passes in the semi-streaming model for general graphs was a significant open problem until it was solved by Fischer–Mitrović–Uitto [FMU22].
@@
Since the matroid intersection problem is a well-known generalization of the bipartite matching problem, there are also several studies on the matroid intersection problem (and its generalizations) in the streaming setting in the literature [CGQ15, FKK18, GJS23, HS24]. For the matroid intersection problem, in the semi-streaming model, we consider algorithms such that the memory usage is $O((r_{1}+r_{2})\textrm{polylog}(r_{1}+r_{2}))$, where $r_{1}$ and $r_{2}$ denote the ranks of matroids $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, respectively. We note that this memory requirement is natural when we formulate the bipartite matching problem as the matroid intersection problem. In this work, we focus on constant-passes ($\text{poly}(1/\varepsilon)$-passes) semi-streaming algorithms for the matroid intersection problem, particularly given the significance of constant-passes semi-streaming algorithms for the matching problem. [^5] Based on the results from Assadi [Ass24] and Quanrud [Qua24a], we believe that it is able to develop a semi-streaming $(1-\varepsilon)$-approximation algorithm with similar space usage. However, such an algorithm would require $O(\log(n)\cdot\text{poly}(1/\varepsilon))$ passes. In the context of research on streaming matching algorithms, the distinctions between $O(\text{poly}(1/\varepsilon))$ passes and $O(\log(n)\cdot\text{poly}(1/\varepsilon))$ are well recognized and crucial; see [Ass24, Table 1 in the arXiv version]. We also believe that the result by Blikstad–Tu [BT24] is unlikely to immediately lead to constant-passes semi-streaming algorithm for the matroid intersection problem. This is because, while Blikstad–Tu used some idea from the constant-passes semi-streaming algorithm of Assadi-Liu-Tarjan [ALT21] for the bipartite matching problem, Blikstad–Tu relies on the technique from Quanrud [Qua24a]. Quanrud’s algorithm used some idea from the $O(\log(n)\cdot\text{poly}(1/\varepsilon))$-passes semi-streaming algorithm by Assadi [Ass24]. Recently, Huang–Sellier [HS24] presented a $(2/3-\varepsilon)$-approximation semi-streaming algorithm for the matroid intersection problem in the random-order streaming model. [^6] We note that their algorithm requires too many queries to compute the density-based decomposition. Thus, their algorithm does not imply a nearly-linear-independence-query $(2/3-\varepsilon)$-approximation algorithm.
@@
Theorem 3.
Given two matroids $\mathcal{M}_{1}=(V,\mathcal{I}_{1})$ and $\mathcal{M}_{2}=(V,\mathcal{I}_{2})$ on the same ground set $V$, for any $\varepsilon>0$, there is a deterministic semi-streaming algorithm that finds a common independent set $S\in\mathcal{I}_{1}\cup\mathcal{I}_{2}$ with $|S|\geq(2/3-\varepsilon)r$, using $O(\frac{r_{1}+r_{2}}{\varepsilon})$ memory and $O(\frac{1}{\varepsilon})$ passes.
Our algorithm in Theorem 3 is a generalization of the bipartite matching semi-streaming algorithm of Feigenbaum et al. [FKM+05].
@@ 2.2 Exchange Graph and Augmenting Path
Definition 4 (Exchange Graph).
@@
Lemma 5 (Shortest Augmenting Path; see [Sch03, Theorem 41.2]).
@@
Cunningham’s [Cun86] matroid intersection algorithm and recent fast matroid intersection algorithms [CQ16, Ngu19, CLS+19, BvdBMN21, Bli21] rely on the following lemma.
Lemma 6 (from [Cun86, Corollary 2.2]).
@@
Lemma 7 (Binary Search Technique from [Ngu19, CLS+19]).
@@
Let FindExchange$(\mathcal{M},S,v,T)$ be the procedure that implements Lemma 7.
Let $D_{i}$ be the set of elements $v\in V$ such that the distance from $s$ to $v$ is exactly $i$ in the exchange graph $G(S)$. Chakrabarty et al. showed that the distance layers $D_{i}$ can be computed efficiently by using the binary search technique; see [CLS+19, Algorithm 6 and Lemma 19]. Their algorithm requires $O(nr+r^{2}\log r)$ independence oracle queries to compute all distance layers $D_{1},D_{2},\ldots,D_{2r+1}$; see [CLS+19, Remark just after Proof of Theorem 18]. In our matroid intersection algorithm, we only need to compute $D_{1}$, $D_{2}$, and $D_{3}$. Their algorithm can compute $D_{1}$, $D_{2}$, and $D_{3}$ using $O(n+r\log r)$ independence oracle queries. See GetDistance (Algorithm 1) for the pseudocode of the algorithm.
@@ 12       $D_{3}\leftarrow D_{3}+v$
Lemma 8 (follows from [CLS+19, Lemma 19]).
@@
For completeness, we give a proof of Lemma 8.
Proof.
The procedure GetDistance (Algorithm 1) simply performs a breadth first search in the exchange graph $G(S)$. Thus, the procedure GetDistance correctly computes the distance layers $D_{1}$, $D_{2}$, and $D_{3}$. Here, each element $u\in D_{2}$ is found by FindExchange only once. Thus, the number of FindExchange calls that do not output $\emptyset$ is $|D_{2}|\leq r$, and the number of FindExchange calls that output $\emptyset$ is $|D_{1}|\leq n$. Hence, by Lemma 7, the number of independence oracle queries used in Line 1 is $O(n+r\log r)$. In addition, the number of independence oracle queries used in Lines 1 and 1 is $O(n)$, which completes the proof. ∎
@@ 3 Augmenting Sets Technique
In this section, we recall the definition and properties of augmenting sets, which were introduced as a generalization of augmenting paths by Chakrabarty–Lee–Sidford–Singla–Wong [CLS+19]. The augmenting sets technique plays a crucial role in the $(1-\varepsilon)$-approximation algorithms of Chakrabarty et al. [CLS+19] and Blikstad [Bli21].
@@
Definition 9 (Augmenting Sets from [CLS+19, Definition 24]).
@@
Theorem 10 (from [CLS+19, Theorem 25]).
@@
Definition 11 (Maximal Augmenting Sets from [CLS+19, Definitions 31 and 35]).
@@
Theorem 12 (from [CLS+19, Theorem 36]).
@@
Definition 13 (Partial Augmenting Sets from [CLS+19, Definition 37]).
@@
Lemma 14 (from [CLS+19, Lemma 38]).
@@
Chakrabarty et al. [CLS+19] presented an algorithm to find a maximal augmenting set in the exchange graph $G(S)$, where shortest augmenting paths have length $2(\ell+1)$, using $O(n^{1.5}\sqrt{\ell\log r})$ independence oracle queries; see [CLS+19, Proof of Theorem 48 and Proof of Theorem 21]. Blikstad [Bli21] improved the query complexity and presented an algorithm to find a maximal augmenting set using $O(n\sqrt{r\log r})$ independence oracle queries; see [Bli21, Lemma 35 and Proof of Theorem 1].
In Blikstad’s algorithm for finding a maximal augmenting set, we keep track of a partial augmenting set and iteratively update it to closer to a maximal augmenting set. To obtain a fast algorithm, Blikstad combines two algorithms Refine [Bli21, Algorithm 4] and RefinePath [Bli21, Algorithm 5]; see [Bli21, Algorithm 6 and Lemma 35]. Let $p\in[1,r]$ be the parameter that controls the trade-off of the two algorithms. He first applies Refine to find a partial augmenting set that is close enough to a maximal augmenting set, which uses $O(nr/p)$ independence oracle queries. Then, he applies RefinePath until the partial augmenting set becomes a maximal augmenting set, which uses $O(np\log r)$ independence oracle queries.
@@
Lemma 15 (follows from [Bli21, Lemma 35]).
@@
To obtain Lemma 15, we apply Refine [Bli21, Algorithm 4] (see also Algorithm 5) until $|B_{1}|-|B_{2}|\leq p$, but at least once. The procedure Refine makes the partial augmenting set $\Phi=(B_{1},A_{1},B_{2})$ close to a maximal augmenting set. To become $|B_{1}|-|B_{2}|\leq p$, we need to apply Refine $O(|S|/p+1)$ times; see [Bli21, Proof of Lemma 35]. Since each call of Refine uses $O(n)$ independence oracle queries (see [Bli21, Lemma 29]), the total number of independence oracle queries is $O(nr/p)$.
Remark 16.
Lemma 15 is not explicitly stated in Blikstad’s [Bli21] paper. By the property of a partial augmenting set and the argument in [Bli21, Proof of Lemma 35], it is clear that the conditions $(i)$ and $(ii)$ are satisfied. In particular, here we show that the condition $(iii)$ is also satisfied. In the algorithm in [Bli21, Algorithm 6 and Lemma 35] for finding a maximal augmenting set, after applying Refine until $|B_{1}|-|B_{2}|\leq p$, we apply RefinePath [Bli21, Algorithm 5] until the partial augmenting set becomes a maximal augmenting set. While RefinePath modifies $B_{1}$, it does not increase $|B_{1}|$; see [Bli21, Proof of Lemma 35]. Thus, the partial augmenting set after applying RefinePath is a maximal augmenting set of width at most $|B_{1}|$. Here, let this maximal augmenting set be $\tilde{\Pi}$.

### v3

@@ Abstract
keywords:
Matroid intersection, approximation algorithm, streaming algorithm
@@ 1 Introduction
The matroid intersection problem is one of the most fundamental problems in combinatorial optimization. In this problem, we are given two matroids $\mathcal{M}_{1}=(V,\mathcal{I}_{1})$ and $\mathcal{M}_{2}=(V,\mathcal{I}_{2})$ defined on the same ground set $V$ of $n$ elements, and the objective is to find a common independent set $S\in\mathcal{I}_{1}\cap\mathcal{I}_{2}$ of largest possible cardinality, denoted by $r$. This problem generalizes many important combinatorial optimization problems such as bipartite matching, packing spanning trees, and arborescences in directed graphs. Furthermore, it has also several applications outside of traditional combinatorial optimization such as electrical engineering [51, 55].
@@
Both the recent $(1-\varepsilon)$-approximation algorithms by Quanrud [53] and Blikstad–Tu [14] are randomized. Therefore, it remains an open question whether a deterministic $\tilde{O}_{\varepsilon}(n)$ independence-query $(1-\varepsilon)$-approximation algorithm can be achieved for almost the entire range of $r$ [^4] Blikstad–Tu [14] presented a deterministic $(1-\varepsilon)$-approximation algorithm using a strictly linear number of independence oracle queries only when $r=\Theta(n)$ (and $\varepsilon$ is constant). Then, we consider a deterministic matroid intersection algorithm with only a nearly linear number of independence oracle queries. It is well-known that any maximal common independent set is at least half the size of a maximum common independent set (see e.g., [47, Proposition 13.26]). Thus, the natural greedy algorithm—pick an element whenever possible—can be regarded as a deterministic $1/2$-approximation algorithm using a linear number of independence oracle queries. In fact, even beating the trivial $1/2$-approximation ratio for deterministic nearly-linear-independence-query algorithms remains an open problem for almost the entire range of $r$44footnotemark: 4
@@
Given two matroids $\mathcal{M}_{1}=(V,\mathcal{I}_{1})$ and $\mathcal{M}_{2}=(V,\mathcal{I}_{2})$ on the same ground set $V$, for any $\varepsilon>0$, there is a deterministic algorithm that finds a common independent set $S\in\mathcal{I}_{1}\cap\mathcal{I}_{2}$ with $|S|\geq(2/3-\varepsilon)r$, using $O(\frac{n}{\varepsilon}+r\log r)$ independence oracle queries.
@@
Our algorithm uses only a strictly linear number of independence oracle queries when $r=O(n/\log n)$ (and $\varepsilon$ is constant). Prior to our work, Guruganesh–Singla [34] presented a randomized $O(n)$-independence-query algorithm that achieves $(1/2+\delta)$-approximation for some small constant $\delta>0$, which was the best approximation ratio for (possibly randomized) algorithms using only a strictly linear number of independence oracle queries.44footnotemark: 4
@@
We note that the time complexity of our algorithm is dominated by the independence oracle queries. That is, our algorithm in Theorem 1.1 has time complexity $O\left((\frac{n}{\varepsilon}+r\log r)\cdot\mathcal{T}_{\textrm{ind}}\right)$, where $\mathcal{T}_{\textrm{ind}}$ denotes the maximum running time of a single independence oracle query.
@@
Our idea is very simple: we apply a recent $\tilde{O}(n\sqrt{r}/\varepsilon)$-independence-query $(1-\varepsilon)$-approximation algorithm of Blikstad [11], but terminate it before completion. We observe that a $(2/3-\varepsilon)$-approximate solution can be obtained by terminating Blikstad’s algorithm early. As such, our algorithm does not introduce technical novelty, but we believe that it enhances the understanding of algorithms for the matroid intersection problem.
@@
We also implement our algorithm using a rank oracle, which takes as input a subset $S\subseteq V$ and outputs the size of the maximum cardinality independent subset of $S$. Several recent studies have considered fast matroid intersection algorithms in the rank oracle model [49, 18, 60]. We note that the rank oracle is more powerful than the independence oracle, since a single query to the rank oracle can determine whether a given set is independent or not.
@@
Chakrabarty–Lee–Sidford–Singla–Wong [18] presented an $\tilde{O}(n\sqrt{r})$-rank-query exact algorithm and an $O(\frac{n}{\varepsilon}\log n)$-rank-query $(1-\varepsilon)$-approximation algorithm. Their $(1-\varepsilon)$-approximation algorithm requires nearly linear rank oracle queries. Our second result is to obtain a $(2/3-\varepsilon)$-approximation algorithm using a strictly linear number of rank oracle queries.
Theorem 1.2.
@@
Given two matroids $\mathcal{M}_{1}=(V,\mathcal{I}_{1})$ and $\mathcal{M}_{2}=(V,\mathcal{I}_{2})$ on the same ground set $V$, for any $\varepsilon>0$, there is a deterministic algorithm that finds a common independent set $S\in\mathcal{I}_{1}\cap\mathcal{I}_{2}$ with $|S|\geq(2/3-\varepsilon)r$, using $O(\frac{n}{\varepsilon})$ rank oracle queries.
@@
We note that the time complexity of our algorithm is dominated by the rank oracle queries. That is, our algorithm in Theorem 1.2 has time complexity $O\left(\frac{n}{\varepsilon}\cdot\mathcal{T}_{\textrm{rank}}\right)$, where $\mathcal{T}_{\textrm{rank}}$ denotes the maximum running time of a single rank oracle query.
@@
Surprisingly, our algorithm can also be implemented in the semi-streaming model of computation. The streaming model for graph problems was initiated by Feigenbaum–Kannan–McGregor–Suri–Zhang [27], which is called the semi-streaming model. In this model, an algorithm computes over a graph stream using a limited space of $\tilde{O}(n)$ bits, where $n$ is the number of vertices. The maximum matching problem is one of the most studied problems in the graph streaming setting [27, 50, 25, 1, 33, 24, 45, 2, 42, 26, 41, 3, 44, 59, 31, 8, 6, 7, 30, 40, 10, 46, 29, 5]. The first of these studies was a $(2/3-\varepsilon)$-approximation algorithm for the bipartite maximum matching problem by Feigenbaum et al. [27], which uses $\tilde{O}(n)$ bits of space and $O(\log(1/\varepsilon)/\varepsilon)$ passes. In particular, whether a $(1-\varepsilon)$ approximation can be achieved in $\text{poly}(1/\varepsilon)$ passes in the semi-streaming model for general graphs was a significant open problem until it was solved by Fischer–Mitrović–Uitto [30].
@@
Since the matroid intersection problem is a well-known generalization of the bipartite matching problem, there are also several studies on the matroid intersection problem (and its generalizations) in the streaming setting in the literature [19, 28, 32, 37, 38, 39, 17]. For the matroid intersection problem, in the semi-streaming model, we consider algorithms such that the memory usage is $O((r_{1}+r_{2})\textrm{polylog}(r_{1}+r_{2}))$, where $r_{1}$ and $r_{2}$ denote the ranks of matroids $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, respectively. We note that this memory requirement is natural when we formulate the bipartite matching problem as the matroid intersection problem. In this work, we focus on constant-passes ($\text{poly}(1/\varepsilon)$-passes) semi-streaming algorithms for the matroid intersection problem, particularly given the significance of constant-passes semi-streaming algorithms for the matching problem. [^5] Based on the results from Assadi [5] and Quanrud [53], we believe that it is able to develop a semi-streaming $(1-\varepsilon)$-approximation algorithm with similar space usage. However, such an algorithm would require $O(\log(n)\cdot\text{poly}(1/\varepsilon))$ passes. In the context of research on streaming matching algorithms, the distinctions between $O(\text{poly}(1/\varepsilon))$ passes and $O(\log(n)\cdot\text{poly}(1/\varepsilon))$ are well recognized and crucial; see [5, Table 1 in the arXiv version]. We also believe that the result by Blikstad–Tu [14] is unlikely to immediately lead to constant-passes semi-streaming algorithm for the matroid intersection problem. This is because, while Blikstad–Tu used some idea from the constant-passes semi-streaming algorithm of Assadi-Liu-Tarjan [8] for the bipartite matching problem, Blikstad–Tu relies on the technique from Quanrud [53]. Quanrud’s algorithm used some idea from the $O(\log(n)\cdot\text{poly}(1/\varepsilon))$-passes semi-streaming algorithm by Assadi [5]. Recently, Huang–Sellier [37] presented a $(2/3-\varepsilon)$-approximation semi-streaming algorithm for the matroid intersection problem in the random-order streaming model. [^6] We note that their algorithm requires too many queries to compute the density-based decomposition. Thus, their algorithm does not imply a nearly-linear-independence-query $(2/3-\varepsilon)$-approximation algorithm.
@@
Theorem 1.3.
Given two matroids $\mathcal{M}_{1}=(V,\mathcal{I}_{1})$ and $\mathcal{M}_{2}=(V,\mathcal{I}_{2})$ on the same ground set $V$, for any $\varepsilon>0$, there is a deterministic semi-streaming algorithm that finds a common independent set $S\in\mathcal{I}_{1}\cap\mathcal{I}_{2}$ with $|S|\geq(2/3-\varepsilon)r$, using $O(\frac{r_{1}+r_{2}}{\varepsilon})$ memory and $O(\frac{1}{\varepsilon})$ passes.
Our algorithm in Theorem 1.3 is a generalization of the bipartite matching semi-streaming algorithm of Feigenbaum et al. [27].
@@ 2.2 Exchange Graph and Augmenting Path
Definition 2.1 (Exchange Graph).
@@
Lemma 2.2 (Shortest Augmenting Path; see [56, Theorem 41.2]).
@@
Cunningham’s [21] matroid intersection algorithm and recent fast matroid intersection algorithms [20, 52, 18, 15, 11] rely on the following lemma.
Lemma 2.3 (from [21, Corollary 2.2]).
@@
Lemma 2.4 (Binary Search Technique from [52, 18]).
@@
Let FindExchange$(\mathcal{M},S,v,T)$ be the procedure that implements Lemma 2.4.
Let $D_{i}$ be the set of elements $v\in V$ such that the distance from $s$ to $v$ is exactly $i$ in the exchange graph $G(S)$. Chakrabarty et al. [18] showed that the distance layers $D_{i}$ can be computed efficiently by using the binary search technique. In their algorithm, they compute odd and even layers separately because FindExchange cannot find both incoming and outgoing edges due to restrictions on the allowed queries; see [18, Algorithm 6 and Lemma 19]. Their algorithm requires $O(nr+r^{2}\log r)$ independence oracle queries to compute all distance layers $D_{1},D_{2},\ldots,D_{2r+1}$; see [18, Remark just after Proof of Theorem 18]. In our matroid intersection algorithm, we only need to compute $D_{1}$, $D_{2}$, and $D_{3}$. Their algorithm can compute $D_{1}$, $D_{2}$, and $D_{3}$ using $O(n+r\log r)$ independence oracle queries. See GetDistance (Algorithm 1) for the pseudocode of the algorithm.
@@ 12       $D_{3}\leftarrow D_{3}+v$
Lemma 2.5 (follows from [18, Lemma 19]).
@@
For completeness, we give a proof of Lemma 2.5.
Proof 2.6.
The procedure GetDistance (Algorithm 1) simply performs a breadth first search in the exchange graph $G(S)$. Thus, the procedure GetDistance correctly computes the distance layers $D_{1}$, $D_{2}$, and $D_{3}$. Here, each element $u\in D_{2}$ is found by FindExchange only once. Thus, the number of FindExchange calls that do not output $\emptyset$ is $|D_{2}|\leq r$, and the number of FindExchange calls that output $\emptyset$ is $|D_{1}|\leq n$. Hence, by Lemma 2.4, the number of independence oracle queries used in Line 1 is $O(n+r\log r)$. In addition, the number of independence oracle queries used in Lines 1 and 1 is $O(n)$, which completes the proof.
@@ 3 Augmenting Sets Technique
In this section, we recall the definition and properties of augmenting sets, which were introduced as a generalization of augmenting paths by Chakrabarty–Lee–Sidford–Singla–Wong [18]. The augmenting sets technique plays a crucial role in the $(1-\varepsilon)$-approximation algorithms of Chakrabarty et al. [18] and Blikstad [11].
@@
Definition 3.1 (Augmenting Sets from [18, Definition 24]).
@@
Theorem 3.2 (from [18, Theorem 25]).
@@
Definition 3.3 (Maximal Augmenting Sets from [18, Definitions 31 and 35]).
@@
Theorem 3.4 (from [18, Theorem 36]).
@@
Definition 3.5 (Partial Augmenting Sets from [18, Definition 37]).
@@
Lemma 3.6 (from [18, Lemma 38]).
@@
Chakrabarty et al. [18] presented an algorithm to find a maximal augmenting set in the exchange graph $G(S)$, where shortest augmenting paths have length $2(\ell+1)$, using $O(n^{1.5}\sqrt{\ell\log r})$ independence oracle queries; see [18, Proof of Theorem 48 and Proof of Theorem 21]. Blikstad [11] improved the query complexity and presented an algorithm to find a maximal augmenting set using $O(n\sqrt{r\log r})$ independence oracle queries; see [11, Lemma 35 and Proof of Theorem 1].
In Blikstad’s algorithm for finding a maximal augmenting set, we keep track of a partial augmenting set and iteratively update it to closer to a maximal augmenting set. To obtain a fast algorithm, Blikstad combines two algorithms Refine [11, Algorithm 4] and RefinePath [11, Algorithm 5]; see [11, Algorithm 6 and Lemma 35]. Let $p\in[1,r]$ be the parameter that controls the trade-off of the two algorithms. He first applies Refine to find a partial augmenting set that is close enough to a maximal augmenting set, which uses $O(nr/p)$ independence oracle queries. Then, he applies RefinePath until the partial augmenting set becomes a maximal augmenting set, which uses $O(np\log r)$ independence oracle queries.
@@
Lemma 3.7 (follows from [11, Lemma 35]).
@@
To obtain Lemma 3.7, we apply Refine [11, Algorithm 4] (see also Algorithm 5) until $|B_{1}|-|B_{2}|\leq p$, but at least once. The procedure Refine makes the partial augmenting set $\Phi=(B_{1},A_{1},B_{2})$ close to a maximal augmenting set. To become $|B_{1}|-|B_{2}|\leq p$, we need to apply Refine $O(|S|/p+1)$ times; see [11, Proof of Lemma 35]. Since each call of Refine uses $O(n)$ independence oracle queries (see [11, Lemma 29]), the total number of independence oracle queries is $O(nr/p)$.
Remark 3.8.
Lemma 3.7 is not explicitly stated in Blikstad’s [11] paper. By the property of a partial augmenting set and the argument in [11, Proof of Lemma 35], it is clear that the conditions $(i)$ and $(ii)$ are satisfied. In particular, here we show that the condition $(iii)$ is also satisfied. In the algorithm in [11, Algorithm 6 and Lemma 35] for finding a maximal augmenting set, after applying Refine until $|B_{1}|-|B_{2}|\leq p$, we apply RefinePath [11, Algorithm 5] until the partial augmenting set becomes a maximal augmenting set. While RefinePath modifies $B_{1}$, it does not increase $|B_{1}|$; see [11, the second paragraph of the proof of Lemma 35]. Thus, the partial augmenting set after applying RefinePath is a maximal augmenting set of width at most $|B_{1}|$. Here, let this maximal augmenting set be $\tilde{\Pi}$.